### v1

# Mieko Ochi

Abstract
Mitotic figures are classified into typical and atypical variants, with atypical counts correlating strongly with tumor aggressiveness. Accurate differentiation is therefore essential for patient prognostication and resource allocation, yet remains challenging even for expert pathologists. Here, we leveraged Pathology Foundation Models (PFMs) pre-trained on large histopathology datasets and applied parameter-efficient fine-tuning via low-rank adaptation. During training, we employ a fisheye transform to emphasize mitoses and Fourier Domain Adaptation using ImageNet target images. Finally, we ensembled multiple PFMs to integrate complementary morphological insights, achieving a high balanced accuracy on the Preliminary Evaluation Phase dataset.
keywords:
ensemble | foundation model | fisheye | fourier domain adaptation | mitosis
{corrauthor}
be_yuan@med.jrc.or.jp
Introduction
Hematoxylin and eosin–stained mitotic figure (MF) counts are essential for tumor evaluation, serving both as standalone and component grades in malignancy assessment (1). Mitotic figures are broadly classified into typical and atypical variants, with atypical forms—characterized by dysregulated chromatin aggregation and reflecting genomic instabilities such as chromosomal instability and aneuploidy—demonstrating independent prognostic value in cancers like breast carcinoma (2, 3). However, manual enumeration and discrimination of MF variants is time-consuming and subject to substantial inter-observer variability.
To address these challenges, we present a two-stage framework for automated MF classification in the MIDOG2025 Track 2 challenge (4). First, we perform parameter-efficient fine-tuning of multiple Pathology Foundation Models (PFMs) using low-rank adaptation (LoRA) (5). Training incorporates fisheye augmentation to emphasize central mitoses (6) and Fourier Domain Adaptation (FDA) for unsupervised style transfer with ImageNet images (7). We further enhance domain generalization by augmenting the MIDOG2025 set with an external labeled MF dataset (8). Second, we ensemble the adapted PFMs to integrate complementary morphological insights into a unified classification decision (9). Our method achieves high balanced accuracy on validation splits, underscoring its potential for reliable, automated MF analysis.
Material and Methods
0.1 Setting of the Training and Validation Datasets
We trained all models using four publicly available datasets: the AMi-Br dataset (10), the MIDOG 2025 Atypical Training Set (11), the OMG-Octo Atypical dataset (12), and the dataset from Mostafa et al. (13). The AMi-Br and MIDOG 2025 Atypical Training Set images were randomly split into training and validation subsets at a 4:1 ratio, and the model achieving the highest accuracy on the validation set was selected for final evaluation. Both the OMG-Octo Atypical and Mostafa et al. datasets were used in their entirety for training.
In preprocessing, each input image was resized to 224 × 224 pixels while preserving its aspect ratio, and then underwent random brightness and contrast adjustments, followed by a random rotation. Next, an optical fisheye distortion was applied with the distortion coefficient sampled uniformly from –0.9 to 0.9. FDA was applied with a probability of $p=0.5$, each time using a randomly selected target image drawn from a pool of 50,000 unlabeled ImageNet images.
For fine-tuning, we used the Adam optimizer with a learning rate of $1\times 10^{-4}$ and a batch size of 32 for all models. Training was performed for 50 epochs with early stopping. Additionally, we employed a WeightedRandomSampler with sampling weights of $1:0.15:0.15$ for the AMi-Br dataset + the MIDOG 2025 Atypical Training Set, OMG-Octo Atypical), and the Mostafa et al. dataset, respectively.
0.2 LoRA for the Foundation Models and Ensemble Learning
Based on the results reported by Banerjee et al. in their LoRA fine-tuning of foundation models on the AMi-Br dataset (10), we selected three PFMs — UNI (14), Virchow (15), and Virchow2 (16) — for our own fine-tuning experiments. Furthermore, following Haotian et al.’s method (17) of applying LoRA to the query (Q) and value (V) projection matrices in the multi-head self-attention (MHSA) module, we introduced two low-rank matrices $A_{Q}\in\mathbb{R}^{d\times r}$ and $B_{Q}\in\mathbb{R}^{r\times k}$, so that the update to the frozen query weight $W_{Q}\in\mathbb{R}^{d\times k}$ can be factorized as
$\Delta W_{Q}$ $=A_{Q}\,B_{Q},$ (1)
$W_{Q}$ $=W_{0}+\Delta W_{Q}=W_{0}+A_{Q}\,B_{Q}\,.$ (2)
Only $A_{Q}$ and $B_{Q}$ are learned during fine-tuning, while the original weight matrix $W_{0}$ remains fixed.
The foundation models’ divergent pretraining datasets lead to substantial variations in feature extraction and diagnostic decision-making. To balance model diversity with clinical reliability, we develop a weighted ensemble framework that aggregates predictions from $N$ fine-tuned models.
Each base model $M_{i}$ produces a probability vector over the $C$ (=2) classes:
$P_{i}(x)=\bigl{[}p_{i}^{(1)}(x),\dots,p_{i}^{(C)}(x)\bigr{]},\quad\sum_{c=1}^{C}p_{i}^{(c)}(x)=1.$ (3)
We then learn nonnegative weights $w_{i}$ on a validation set $\mathcal{D}_{\rm val}$ by maximizing diagnostic accuracy. Traditional overall accuracy maximization (Equation 4) may sacrifice performance on minority classes when class imbalance persists in the validation set. Therefore, we introduce a new objective function that directly maximizes balanced accuracy.
$\mathbf{w}^{*}=\arg\max_{\mathbf{w}}$ $\frac{1}{C}\sum_{c=1}^{C}\frac{1}{|D_{\mathrm{val}}^{c}|}$ (4)
$\quad\times\sum_{(x,y)\in D_{\mathrm{val}}^{c}}\mathbf{1}\Bigl{[}y=\arg\max_{c^{\prime}}\sum_{i=1}^{N}w_{i}\,p_{i}^{(c^{\prime})}(x)\Bigr{]}$
$\mathrm{s.t.}$ $w_{i}\geq 0,\quad\sum_{i=1}^{N}w_{i}=1.$
Here, $C$ denotes the number of classes and $D_{\mathrm{val}}^{c}$ denotes the set of validation samples with label $c$. By equally weighting the accuracy of each class, we mitigate performance degradation on minority classes.
Finally, the ensemble prediction is
$P_{\rm final}^{(c)}(x)=\sum_{i=1}^{N}w_{i}^{*}\,p_{i}^{(c)}(x),\quad\hat{y}=\arg\max_{c\in\{1,\dots,C\}}P_{\rm final}^{(c)}(x).$ (5)
Results
As shown in Table 1, we conducted an ablation study using the UNI models on our validation dataset to evaluate the effect of fisheye and FDA augmentations during training. The highest balanced accuracy was achieved when fisheye augmentation was applied without FDA (85.457%), followed by the combination of both fisheye and FDA (83.993%).
Table 1: Ablation of fisheye and FDA augmentation of UNI models on our validation dataset.
Fisheye FDA Balanced Acc (%)
83.155
85.457
83.519
83.993
Furthermore, under the same training augmentation conditions, we compared the performance of each PFM as well as their ensemble on the validation dataset (Table 2). The best balanced accuracy was achieved by Virchow2, and ensembling all PFMs resulted in an improvement of approximately 5% in balanced accuracy.
Table 2: Performance comparison on our validation dataset.
Category Model Balanced Acc (%)
Single Model UNI 85.457
Virchow 86.035
Virchow2 87.590
Ensemble Model UNI + Virchow + Virchow2 93.568
Based on these results, in the Preliminary Evaluation Phase, we submitted an ensemble model trained under two conditions: fisheye only (Fisheye Only) and fisheye combined with FDA augmentation (Fisheye + FDA), and compared their performance. As Table 3 shows, while the overall balanced accuracy was higher with Fisheye + FDA, in domain_0, its performance dropped by approximately 23% compared to Fisheye Only. In other words, Fisheye Only exhibited stable performance across all domains, whereas Fisheye + FDA achieved strong performance in some domains and showed greater variability across domains.
Table 3: Comparison of balanced accuracy (%) between the submitted models in the Preliminary Evaluation Phase.
Fisheye Only Fisheye + FDA
domain_0 76.563 78.125
domain_1 83.843 85.188
domain_2 88.764 91.011
domain_3 94.444 95.833
overall balanced accuracy 86.803 88.371
Discussion
We based our approach on the first-place solution by Haotian et al (17) in the Pap Smear Cell Classification Challenge, combining parameter–efficient LoRA fine-tuning of individual PFMs with a subsequent ensemble learning stage. Banerjee et al. (18) demonstrated that, among the diverse PFMs evaluated on the AMi-Br dataset (one of the challenge’s benchmarks) (10), Virchow2, UNI, and Virchow achieved the highest balanced accuracies. Accordingly, we selected these three models for per-model adaptation. Each PFM has been pretrained on large collections of human histopathology images acquired under different conditions (institutions, countries, staining protocols, scanners, etc.), resulting in complementary attention patterns over cellular structures. By ensembling their predictions, we exploit these diverse inductive biases to arrive at more accurate final diagnoses.
Because distinguishing normal from atypical mitotic figures requires capturing subtle chromatin texture variations, we further augmented our training set with a fisheye transformation that emphasizes pixels in the central region of each image—where mitotic figures typically reside. Fisheye augmentation has previously been shown to boost single-cell classification accuracy in bladder, lung, and other tissues (6). Moreover, to mitigate domain shifts arising from differences in scanner types and staining protocols, we applied style transfer via FDA using generic natural images. Prior work by Yamashita et al. (19) and Yang et al.(7) has shown that natural image–based style transfer, including FDA, can substantially improve model performance across a variety of histopathological tasks—ranging from tumor gene subtyping to mitotic figure detection. Our results clearly demonstrate the efficacy of this tailored adaptation strategy for histopathological image analysis and have the potential to reduce pathologists’ workload in clinical practice. In future work, we will explore advanced techniques to further enhance model generalization and validate our approach on larger, more diverse clinical datasets.

### v2

@@
 Here, we leveraged Pathology Foundation Models (PFMs) pre-trained on large histopathology datasets and applied parameter-efficient fine-tuning via low-rank adaptation.
-During training, we employ a fisheye transform to emphasize mitoses and Fourier Domain Adaptation using ImageNet target images.
-Finally, we ensembled multiple PFMs to integrate complementary morphological insights, achieving a high balanced accuracy on the Preliminary Evaluation Phase dataset.
+In addition, we incorporated ConvNeXt V2, a state-of-the-art convolutional neural network architecture, to complement PFMs.
+During training, we employed a fisheye transform to emphasize mitoses and Fourier Domain Adaptation using ImageNet target images.
+Finally, we ensembled multiple PFMs to integrate complementary morphological insights, achieving competitive balanced accuracy on the Preliminary Evaluation Phase dataset.
 keywords:
@@
 {corrauthor}
-be_yuan@med.jrc.or.jp
+noonecanpass.and@gmail.com, be_yuan@med.jrc.or.jp
 Introduction
@@
 Mitotic figures are broadly classified into typical and atypical variants, with atypical forms—characterized by dysregulated chromatin aggregation and reflecting genomic instabilities such as chromosomal instability and aneuploidy—demonstrating independent prognostic value in cancers like breast carcinoma (2, 3).
-However, manual enumeration and discrimination of MF variants is time-consuming and subject to substantial inter-observer variability.
+However, manual enumeration and discrimination of MF variants are time-consuming and subject to substantial inter-observer variability.
 To address these challenges, we present a two-stage framework for automated MF classification in the MIDOG2025 Track 2 challenge (4).
-First, we perform parameter-efficient fine-tuning of multiple Pathology Foundation Models (PFMs) using low-rank adaptation (LoRA) (5).
-Training incorporates fisheye augmentation to emphasize central mitoses (6) and Fourier Domain Adaptation (FDA) for unsupervised style transfer with ImageNet images (7).
-We further enhance domain generalization by augmenting the MIDOG2025 set with an external labeled MF dataset (8).
-Second, we ensemble the adapted PFMs to integrate complementary morphological insights into a unified classification decision (9).
-Our method achieves high balanced accuracy on validation splits, underscoring its potential for reliable, automated MF analysis.
+First, we performed parameter-efficient fine-tuning of multiple Pathology Foundation Models (PFMs) using low-rank adaptation (LoRA) (5).
+Training incorporated fisheye augmentation to emphasize central mitoses (6) and Fourier Domain Adaptation (FDA) for unsupervised style transfer with ImageNet images (7).
+We further enhanced domain generalization by augmenting the MIDOG2025 set with an external labeled MF dataset (8).
+Second, we ensembled the adapted PFMs and ConvNeXt V2 (9) to integrate complementary morphological insights into a unified classification decision (10).
+Our method achieved a high balanced accuracy on validation splits and also demonstrated strong performance on the Preliminary Evaluation Phase dataset, underscoring its potential for reliable, automated MF analysis.
 Material and Methods
@@
 (14).
-The AMi-Br and MIDOG 2025 Atypical Training Set images were randomly split into training and validation subsets at a 4:1 ratio, and the model achieving the highest accuracy on the validation set was selected for final evaluation.
-Both the OMG-Octo Atypical and Mostafa et al.
-datasets were used in their entirety for training.
-In preprocessing, each input image was resized to 224 × 224 pixels while preserving its aspect ratio, and then underwent random brightness and contrast adjustments, followed by a random rotation.
-Next, an optical fisheye distortion was applied with the distortion coefficient sampled uniformly from –0.9 to 0.9.
+The AMi-Br and MIDOG 2025 Atypical Training Set images were randomly split into training and validation subsets at a 4:1 ratio.
+The OMG-Octo Atypical and Mostafa et al.
+datasets were each used in their entirety for training.
+In preprocessing, each input image was resized to 224 × 224 pixels with aspect ratio preserved and padded as necessary, and then underwent random brightness and contrast adjustments, followed by a random rotation.
+Next, an optical fisheye distortion was applied with the distortion coefficient sampled uniformly sampled from –0.9 to 0.9.
 FDA was applied with a probability of $p=0.5$, each time using a randomly selected target image drawn from a pool of 50,000 unlabeled ImageNet images.
-For fine-tuning, we used the Adam optimizer with a learning rate of $1\times 10^{-4}$ and a batch size of 32 for all models.
+For fine-tuning, we used the Adam optimizer (learning rate $1{\times}10^{-4}$, weight decay $1{\times}10^{-6}$ for ConvNeXt V2 and $1\times 10^{-4}$ without weight deca for PFMs) and a batch size of 32 for all models.
 Training was performed for 50 epochs with early stopping.
@@
 0.2 LoRA for the Foundation Models and Ensemble Learning
-Based on the results reported by Banerjee et al.
-in their LoRA fine-tuning of foundation models on the AMi-Br dataset (10), we selected three PFMs — UNI (14), Virchow (15), and Virchow2 (16) — for our own fine-tuning experiments.
-Furthermore, following Haotian et al.’s method (17) of applying LoRA to the query (Q) and value (V) projection matrices in the multi-head self-attention (MHSA) module, we introduced two low-rank matrices $A_{Q}\in\mathbb{R}^{d\times r}$ and $B_{Q}\in\mathbb{R}^{r\times k}$, so that the update to the frozen query weight $W_{Q}\in\mathbb{R}^{d\times k}$ can be factorized as
+We fine-tuned ImageNet-pretrained ConvNeXt V2 (base, fcmae_ft_in22k_in1k) with a binary classification head.
+For PFMs, based on the results reported by Banerjee et al.
+from LoRA fine-tuning of foundation models on the AMi-Br dataset (11), we selected three PFMs — UNI (15), Virchow (16), and Virchow2 (17) — for our own LoRA fine-tuning experiments.
+Furthermore, following Haotian et al.’s method (18) of applying LoRA to the query (Q) and value (V) projection matrices in the multi-head self-attention (MHSA) module, we introduced two low-rank matrices $A_{Q}\in\mathbb{R}^{d\times r}$ and $B_{Q}\in\mathbb{R}^{r\times k}$, such that the update to the frozen query weight $W_{Q}\in\mathbb{R}^{d\times k}$ can be factorized as
 $\Delta W_{Q}$ $=A_{Q}\,B_{Q},$ (1)
@@
 Only $A_{Q}$ and $B_{Q}$ are learned during fine-tuning, while the original weight matrix $W_{0}$ remains fixed.
-The foundation models’ divergent pretraining datasets lead to substantial variations in feature extraction and diagnostic decision-making.
-To balance model diversity with clinical reliability, we develop a weighted ensemble framework that aggregates predictions from $N$ fine-tuned models.
-Each base model $M_{i}$ produces a probability vector over the $C$ (=2) classes:
+The foundation models’ divergent pretraining datasets led to substantial variations in feature extraction and diagnostic decision-making.
+To balance model diversity with clinical reliability, we developed a weighted ensemble framework that aggregates predictions from $N$ fine-tuned models.
+Each base model $M_{i}$ produced a probability vector over the $C$ (=2) classes:
 $P_{i}(x)=\bigl{[}p_{i}^{(1)}(x),\dots,p_{i}^{(C)}(x)\bigr{]},\quad\sum_{c=1}^{C}p_{i}^{(c)}(x)=1.$ (3)
@@
 By equally weighting the accuracy of each class, we mitigate performance degradation on minority classes.
-Finally, the ensemble prediction is
+Finally, the ensemble prediction is given by
 $P_{\rm final}^{(c)}(x)=\sum_{i=1}^{N}w_{i}^{*}\,p_{i}^{(c)}(x),\quad\hat{y}=\arg\max_{c\in\{1,\dots,C\}}P_{\rm final}^{(c)}(x).$ (5)
 Results
-As shown in Table 1, we conducted an ablation study using the UNI models on our validation dataset to evaluate the effect of fisheye and FDA augmentations during training.
+As shown in Table 1, we conducted an ablation study using the UNI model on our validation dataset to assess the impact of fisheye and FDA augmentations during training.
 The highest balanced accuracy was achieved when fisheye augmentation was applied without FDA (85.457%), followed by the combination of both fisheye and FDA (83.993%).
@@
 83.993
-Furthermore, under the same training augmentation conditions, we compared the performance of each PFM as well as their ensemble on the validation dataset (Table 2).
-The best balanced accuracy was achieved by Virchow2, and ensembling all PFMs resulted in an improvement of approximately 5% in balanced accuracy.
+Furthermore, under the same augmentation settings, we compared the performance of each PFM and their ensemble on the validation dataset (Table 2).
+ConvNeXt V2 achieved the highest balanced accuracy among individual models, and ensembling all models yielded an improvement of approximately 10% in balanced accuracy.
 Table 2: Performance comparison on our validation dataset.
@@
 Virchow2 87.590
+ConvNeXt V2 87.745
 Ensemble Model UNI + Virchow + Virchow2 93.568
-Based on these results, in the Preliminary Evaluation Phase, we submitted an ensemble model trained under two conditions: fisheye only (Fisheye Only) and fisheye combined with FDA augmentation (Fisheye + FDA), and compared their performance.
-As Table 3 shows, while the overall balanced accuracy was higher with Fisheye + FDA, in domain_0, its performance dropped by approximately 23% compared to Fisheye Only.
-In other words, Fisheye Only exhibited stable performance across all domains, whereas Fisheye + FDA achieved strong performance in some domains and showed greater variability across domains.
-Table 3: Comparison of balanced accuracy (%) between the submitted models in the Preliminary Evaluation Phase.
-Fisheye Only Fisheye + FDA
-domain_0 76.563 78.125
-domain_1 83.843 85.188
-domain_2 88.764 91.011
-domain_3 94.444 95.833
-overall balanced accuracy 86.803 88.371
+UNI + Virchow + Virchow2 + ConvNeXt V2 97.279
+Based on these results, in the Preliminary Evaluation Phase, we submitted ensemble models of PFMs trained under two conditions: fisheye only (Fisheye Only) and fisheye combined with FDA augmentation (Fisheye + FDA).
+As shown in Table 3, the overall balanced accuracy was higher with Fisheye + FDA, which consistently outperformed Fisheye Only across all domains.
+In other words, Fisheye + FDA exhibited stable performance, whereas Fisheye Only achieved strong results in some domains but showed greater variability overall.
+Finally, we evaluated an ensemble comprising all models, including ConvNeXt V2 trained with Fisheye + FDA augmentation, in the Preliminary Evaluation Phase.
+This ensemble achieved the highest overall balanced accuracy across all configurations.
+Table 3: Comparison of balanced accuracy (%) between the submitted models across domains in the Preliminary Evaluation Phase.
+PFM denotes the ensemble of UNI, Virchow, and Virchow2, while ALL denotes the ensemble of UNI, Virchow, Virchow2, and ConvNeXt V2.
+OBA denotes overall balanced accuracy.
+Fisheye Only (PFM) Fisheye + FDA (PFM) Fisheye + FDA (ALL)
+domain_0 76.563 78.125 95.312
+domain_1 83.843 85.188 83.255
+domain_2 88.764 91.011 91.573
+domain_3 94.444 95.833 94.444
+OBA 86.803 88.371 88.879
 Discussion
@@
 Banerjee et al.
-(18) demonstrated that, among the diverse PFMs evaluated on the AMi-Br dataset (one of the challenge’s benchmarks) (10), Virchow2, UNI, and Virchow achieved the highest balanced accuracies.
+(19) demonstrated that, among the diverse PFMs evaluated on the AMi-Br dataset (one of the challenge’s benchmarks) (11), Virchow2, UNI, and Virchow achieved the high balanced accuracies.
 Accordingly, we selected these three models for per-model adaptation.
@@
 By ensembling their predictions, we exploit these diverse inductive biases to arrive at more accurate final diagnoses.
-Because distinguishing normal from atypical mitotic figures requires capturing subtle chromatin texture variations, we further augmented our training set with a fisheye transformation that emphasizes pixels in the central region of each image—where mitotic figures typically reside.
+In addition, we incorporated ConvNeXt V2, a convolutional neural network (CNN) that achieved state-of-the-art performance across multiple vision benchmarks, including ImageNet recognition, in 2023, into the ensemble.
+Owing to its architectural design, CNNs extract features within local receptive fields of convolutional kernels, thereby capturing local patterns.
+In contrast, Vision Transformers extract features in parallel across the entire image, resulting in global receptive fields and holistic representations (20, 21).
+Given that the target regions of interest in our task—mitotic figures—are confined to small local areas, we hypothesized that CNNs would provide superior accuracy.
+Because distinguishing normal from atypical mitotic figures requires capturing subtle chromatin texture variations, we further augmented our training set with a fisheye transformation that emphasizes the central region of each image—where mitotic figures typically reside.
 Fisheye augmentation has previously been shown to boost single-cell classification accuracy in bladder, lung, and other tissues (6).
-Moreover, to mitigate domain shifts arising from differences in scanner types and staining protocols, we applied style transfer via FDA using generic natural images.
+Moreover, to mitigate domain shifts arising from differences in scanner types and staining protocols, we applied style transfer via FDA using unlabeled natural images.
 Prior work by Yamashita et al.
-(19) and Yang et al.(7) has shown that natural image–based style transfer, including FDA, can substantially improve model performance across a variety of histopathological tasks—ranging from tumor gene subtyping to mitotic figure detection.
-Our results clearly demonstrate the efficacy of this tailored adaptation strategy for histopathological image analysis and have the potential to reduce pathologists’ workload in clinical practice.
+(22) and Yang et al.(7) has shown that natural image–based style transfer, including FDA, have been shown to substantially improve model performance across a variety of histopathological tasks—ranging from tumor gene subtyping to mitotic figure detection.
+Our results demonstrate the efficacy of this tailored adaptation strategy for histopathological image analysis and highlight its potential to reduce pathologists’ workload in clinical practice.
 In future work, we will explore advanced techniques to further enhance model generalization and validate our approach on larger, more diverse clinical datasets.

### v3

@@
 Mitotic figures are broadly classified into typical and atypical variants, with atypical forms—characterized by dysregulated chromatin aggregation and reflecting genomic instabilities such as chromosomal instability and aneuploidy—demonstrating independent prognostic value in cancers like breast carcinoma (2, 3).
-However, manual enumeration and discrimination of MF variants are time-consuming and subject to substantial inter-observer variability.
+However, manual enumeration and discrimination of MF variants are time-consuming and subject to substantial inter‐observer variability.
 To address these challenges, we present a two-stage framework for automated MF classification in the MIDOG2025 Track 2 challenge (4).
@@
 from LoRA fine-tuning of foundation models on the AMi-Br dataset (11), we selected three PFMs — UNI (15), Virchow (16), and Virchow2 (17) — for our own LoRA fine-tuning experiments.
-Furthermore, following Haotian et al.’s method (18) of applying LoRA to the query (Q) and value (V) projection matrices in the multi-head self-attention (MHSA) module, we introduced two low-rank matrices $A_{Q}\in\mathbb{R}^{d\times r}$ and $B_{Q}\in\mathbb{R}^{r\times k}$, such that the update to the frozen query weight $W_{Q}\in\mathbb{R}^{d\times k}$ can be factorized as
+Furthermore, following Haotian et al.’s method (18) of applying LoRA to the query (Q) and value (V) projection matrices in the multi-head self-attention (MHSA) module, we introduced two low‐rank matrices $A_{Q}\in\mathbb{R}^{d\times r}$ and $B_{Q}\in\mathbb{R}^{r\times k}$, such that the update to the frozen query weight $W_{Q}\in\mathbb{R}^{d\times k}$ can be factorized as
 $\Delta W_{Q}$ $=A_{Q}\,B_{Q},$ (1)
 $W_{Q}$ $=W_{0}+\Delta W_{Q}=W_{0}+A_{Q}\,B_{Q}\,.$ (2)
-Only $A_{Q}$ and $B_{Q}$ are learned during fine-tuning, while the original weight matrix $W_{0}$ remains fixed.
+Only $A_{Q}$ and $B_{Q}$ are learned during fine‐tuning, while the original weight matrix $W_{0}$ remains fixed.
 The foundation models’ divergent pretraining datasets led to substantial variations in feature extraction and diagnostic decision-making.
@@
 By equally weighting the accuracy of each class, we mitigate performance degradation on minority classes.
+Equation 4 defines our proposed objective for the ensemble stage (i.e., maximizing balanced accuracy) under simplex constraints on the weights.
+The ensemble weights were not updated jointly with base models.
+Instead, all individual models were fully trained first, and the weights were then optimized post hoc on the validation set using AutoGluon’s weighted ensemble (Caruana-style Ensemble Selection) (19).
+Concretely, we employ a greedy forward selection with replacement over the candidate model library, iteratively adding the model that yields the largest increase in balanced accuracy on out-of-fold/validation predictions, and normalizing the nonnegative weights to sum to 1.
+This procedure is implemented via AutoGluon Tabular (v1.2) and corresponds to its default weighted-ensemble meta-learner.
 Finally, the ensemble prediction is given by
@@
 Discussion
-We based our approach on the first-place solution by Haotian et al (18) in the Pap Smear Cell Classification Challenge, combining parameter–efficient LoRA fine-tuning of individual PFMs with a subsequent ensemble learning stage.
+We based our approach on the first‐place solution by Haotian et al.
+(18) in the Pap Smear Cell Classification Challenge, combining parameter–efficient LoRA fine-tuning of individual PFMs with a subsequent ensemble learning stage.
 Banerjee et al.
@@
 Owing to its architectural design, CNNs extract features within local receptive fields of convolutional kernels, thereby capturing local patterns.
-In contrast, Vision Transformers extract features in parallel across the entire image, resulting in global receptive fields and holistic representations (20, 21).
+In contrast, Vision Transformers extract features in parallel across the entire image, resulting in global receptive fields and holistic representations (21, 22).
 Given that the target regions of interest in our task—mitotic figures—are confined to small local areas, we hypothesized that CNNs would provide superior accuracy.